# Magnetic Skyrmions: from lumps to supercompactons

### v2

@@ 2.3 $\operatorname{O}(3)$ sigma model limit
with $a$ an arbitrary scale of the soliton (size modulus of the $\operatorname{O}(3)$ sigma model lump). We then evaluate the energy functional for small but finite values of $\epsilon$. It turns out that the energy functional (2.35) on this solution (2.43) as a function of $a$, by minimization with respect to $a$, gives rise to an incorrect result. Since we are considering a lump, the falloff of the profile function $\vartheta$ is power-like (see Eq. (2.37)) making the boundary terms in Eqs. (2.40) and (2.41) nonvanishing. It is thus necessary to consider the right energy functional, the one that gives the correct variational problem. The correct choice is either (2.40) or (2.41) without boundary terms. Evaluation of the full energy functional, without the boundary term of the DM, on the lump solution gives
$E(a)=4\pi-8\pi\kappa a+\frac{2^{p}}{p-1}h\pi a^{2}\ ,$ (2.44)
@@
For $p=1$ the solution for the lump (2.43) evaluated on $E_{0}$ diverges, see Eq. (2.44). Introducing a cutoff gives
$E(a)\simeq 4\pi-8\pi\kappa a+2\pi a^{2}h\log\left(1+\frac{\Lambda^{2}}{a^{2}}\right)\ .$ (2.47)
@@
Regularization of the infinity amounts to removing the term $\Lambda$, for which
$a\simeq-\ell\ .$ (2.50)
Unfortunately, this yields a negative lump size which is not physically viable.
@@ 2.4 Restricted model limit
In the case of $\ell$ fixed and $\epsilon\to\infty$ we expect that the stable Skyrmion is essentially given by a stabilization between $E_{1}$ and $E_{0}$, while $E_{2}$ is negligible with respect to the other two terms. Thus the energy functional takes the simplified form333 One may wonder why we restrict the potential to the form $V_{p}$ of Eq. (2.5) and hence do not consider e.g. $V\propto(1-(n^{3})^{2})$. Although such easy-axis potential is often utilized in magnetic systems, they do not allow for the large-$\epsilon$ limit because the ground state becomes inhomogeneous [52, 53] and requires a different analysis than carried out here.
@@ 2.5 Critical coupling
We now consider what happens if we again drop the boundary term of Eq. (2.16) as we had to in the lump case of Sec. 2.3. The energy is thus changed by a total derivative, yielding now
$E$ $=2\pi\int r\mathop{}\!\mathrm{d}r\left[\frac{1}{2}\left(\vartheta^{\prime}-\frac{\sin\vartheta}{r}+\kappa(1-\cos\vartheta)\right)^{2}+\frac{\sin\vartheta}{r}\vartheta^{\prime}+\frac{\kappa}{r}\partial_{r}(r\sin\vartheta(1-\cos\vartheta))\right]$
$=\pi\int r\mathop{}\!\mathrm{d}r\left(\vartheta^{\prime}-\frac{\sin\vartheta}{r}+\kappa(1-\cos\vartheta)\right)^{2}+4\pi Q+4\pi\Omega^{\prime}\ .$ (2.67)
Notice that a total derivative other than the topological charge remains
$\Omega^{\prime}=\frac{\kappa}{2}\int\mathop{}\!\mathrm{d}r\,\partial_{r}(r\sin\vartheta(1-\cos\vartheta))=0\ .$ (2.68)
which however vanishes when evaluated on the BPS solution. This means that dropping the boundary term in Eq. (2.16) leads to a Bogomol’nyi energy (bound)
$E\geq-4\pi\ .$ (2.69)
@@ 2.6 Analytic solution at any coupling
We note that the Bogomol’nyi completion requires us to select the critical coupling $h=\frac{\kappa^{2}}{2}$ enabling us to complete the square in the energy functional. This choice of coupling corresponds to $\epsilon=2$. However, another interesting fact about the Zeeman squared potential ($p=2$) is that inserting the Ansatz $\vartheta=2\arctan\tfrac{a}{r}$ into the full second-order Euler-Lagrange equation gives [19]
@@
which beautifully has the solution $a=\frac{\kappa}{h}=\ell$. This size does not contradict the BPS solution, since the first-order BPS equation required $h=\frac{\kappa^{2}}{2}$ which in the full second-order solution corresponds to the lump size $a=\frac{\kappa}{h}=\frac{2}{\kappa}$. It is interesting, however to insert the exact solution into the energy functional (2.35), which gives
@@
The $p=2$ case is thus applicable to many more physical systems, than just the critically coupled one [19]. Having the analytic solution in hand may be advantageous for further studies, including interaction and scattering etc.
In Sec. (2.3), we found however that the lump-type Ansatz gave the incorrect lump size $a$ by using the energy (2.35), and we should instead drop the boundary term of Eq. (2.40) yielding the energy
$E=2\pi\int r\mathop{}\!\mathrm{d}r\;\frac{4h\kappa^{2}(h^{3}r^{2}+h(1-3hr^{2})\kappa^{2}+\kappa^{4})}{(h^{2}r^{2}+\kappa^{2})^{3}}=4\pi(1-\epsilon)\ .$ (2.72)
@@ 2.7 Numerical solutions
For the general theory, we can solve equation (2.36) using a standard shooting method, beginning from the origin with the initial condition $\vartheta(0)=\pi$. The form of the solution close to the origin is $\vartheta(r)\simeq\pi-\alpha r$. By varying the angular coefficient $c$, we find only one solution in the topological sector $Q=-1$ as already shown in the literature [16, 18, 20, 19, 51, 23, 21].
@@
Figure 4: Topological charge density, $\mathcal{Q}$ for (a) $p=\frac{3}{2}$ and (b) $p=1$ with $\epsilon=0.001$, $0.01$, $0.1$, $1$, $10$, $100$, $1000$. The topological charge densities of the lumps and restricted Skyrmion are shown with dashed black and red curves, respectively. In the $p=1$ case, the size of the lump is selected for illustrative purpose, corresponding to the solutions in Fig. 3.
@@
Figure 5: Energy density for: (a) $p=\frac{3}{2}$ (b) $p=1$ and $\epsilon=0.001,0.01,0.1,1$ with the boundary term of Eq. (2.40) (solid blue lines) and without it (dashed dark-blue lines), and in (c) $p=\frac{3}{2}$ (d) $p=1$ and $\epsilon=1,10,100,1000$ with the boundary term and in (e) $p=\frac{3}{2}$ (f) $p=1$ and $\epsilon=1,10,100,1000$ without the boundary term of Eq. (2.40).
Fig. 5 shows the energy densities for the sigma-model lump limit ($\epsilon\to 0$) in panels (a) and (b) with and without the boundary term for $p=\frac{3}{2}$ and $p=1$, respectively, and for the restricted limit ($\epsilon\to\infty$) in panels (c) and (d) for $p=\frac{3}{2}$ and $p=1$, respectively, with the boundary term and in panels (e) and (f) $p=\frac{3}{2}$ and $p=1$, respectively, without the boundary term of Eq. (2.40). In both the charge density as well as in the energy density, the discontinuity that develops in the $p=1$ case for $\epsilon\to\infty$ is clearly seen in Figs. 4(b) and 5(d), respectively. We note that the energy density is divergent (singular) in the $p=\frac{3}{2}$ case at $\rho=1$ when the boundary term is taken into account, see Fig. 5(c), but is convergent (finite) when the boundary term is dropped, see Fig. 5(e). We also note that the total energy (the integral) is unchanged in the $p=1$ case (see Figs. 5(d) and 5(f)), but the energy density is slightly modified by the presence of the boundary term.
Figure 6: Energy in the lump limit ($\epsilon\to 0$) as a function of $\epsilon$ for $p=\tfrac{3}{2}$ and $p=1$. (a) The blue dashed line corresponds to the analytical energy formula (2.46) for the lump in the limit of small but nonvanishing $\epsilon$ in the case of $p=\frac{3}{2}$. The red dashed line instead is the energy of Eq. (2.51) with the lump size taken from Fig. 3. (b) The energy shown up to larger values of $\epsilon$ revealing the sign change near $21<\epsilon<$ 2. The case $p=\frac{3}{2}$ is shown with (solid blue line) and without the boundary term (dashed blue line) of Eq. (2.40).
We show the total energy of the numerical solutions in the lump limit ($\epsilon\to 0$) for the $p=\frac{3}{2},1$ cases in Fig. 6(a). We notice that the $p=\frac{3}{2}$ case is well described by the energy functional (2.46) with the correct lump size (2.45). This confirms that discarding the boundary term in Eq. (2.40) is appropriate for computing the energies in the lump limit. In the $p=1$ case, we show the regularized energy (2.47) with a dashed red curve and lump size $a$ as chosen in Fig. 3. We show the sign change of the energy in the $p=\frac{3}{2},1$ case in Fig. 6(b) which appears to take place around $21<\epsilon<$ 2, which is consistent with the lump prediction (2.46) which for $p=\frac{3}{2}$ should take place at $\epsilon=\epsilon^{\rm crit}=\sqrt{2}$.
@@ 2.8 Comparison with experiments
We have confirmed that the magnetic Skyrmion model, up to rescaling of energy and length, is essentially a one-parameter model (2.24)-(2.25). This parameter, $\epsilon$, controls the relation between $E_{2}$ and $E_{1}+E_{0}$. In both the large-$\epsilon$ and the small-$\epsilon$ limit, analytic approximations are available, see Secs. 2.3 and 2.4. It may be interesting to explore various phenomenological realizations of chiral magnets and determine the order of magnitude of the parameter $\epsilon$ that can be realized.
@@
where $\operatorname{J}$ is the unit Joule, not to be confused with the interaction strength of the Heisenberg Hamiltonian. Note that in general $\ell$ can change significantly, viz. one order of magnitude. For instance, if the we fix $\ell=100\operatorname{nm}$ and $\lambda\in\mathcal{O}(10^{1})$ nm as before, we find that $\epsilon$ changes one order of magnitude, i.e. $\epsilon\simeq 100$. For materials other than MnSi, e.g. Fe1–xCoxSi or Mn1–xFexGe, $\lambda$ can also be of the order $\mathcal{O}(10^{2})\operatorname{nm}$. Therefore, considering the other two cases where $\lambda$ is fixed as $\lambda\sim 10^{2}\operatorname{nm}$ and $\ell$ takes its maximum and minimum values, we find the relevant ranges for $\epsilon$:
@@
In Secs. 2.5 and 2.6, we saw that in the special physical case of $p=2$, which describes chiral magnets with anisotropy interactions, any spherical magnetic Skyrmion can be represented by the analytical lump solution, regardless of its DM coupling, $\kappa$. However, the potential $V_{p=2}$ corresponds to the sum of a magnetic field and anisotropy (easy-plane) contributions. To realize this potential, the couplings of the anisotropy term must be exactly half of that of the Zeeman term. It turns out to be possible [55]. Furthermore, since the critical coupling regime corresponds to the condition $\forall\leavevmode\nobreak\ \ell,\ \epsilon=2$, which satisfies the inequalities (2.76), we suggest that, in principle, a BPS magnetic Skyrmion could be realized experimentally.
@@ 3 Conclusion
In this paper, we have studied magnetic Skyrmions in chiral magnets with three different potentials, corresponding to $p=1$, $p=\frac{3}{2}$ and $p=2$ in the limits of the control parameter $\epsilon=\frac{\kappa^{2}}{h}=\frac{4D^{2}}{|B|J}$ becoming very large and going to zero. The $p=1$ case is simply the Zeeman potential term and the $p=2$ case is a particular combination of the easy-plane (anisotropy) and the Zeeman potentials. The $p=\frac{3}{2}$ case is probably somewhat more academic. The $\epsilon$ vanishing limit reduces the Skyrmion to a sigma model lump with a lump size that we accurately estimated, except for the case of the Zeeman term ($p=1$). The large-$\epsilon$ limit instead corresponds to the “restricted” magnetic Skyrmions, for which we found the exact analytic solutions as well. In the $p=\frac{3}{2}$ case, the restricted solution is a compacton, similar in nature to those found in the baby-Skyrme [43, 44, 45] and 3D Skyrme models [46]. However, in the $p=1$ case corresponding to the Zeeman term, we find a new type of restricted solution, which we dub a supercompacton, since it becomes discontinuous. We provide numerical solutions showing that they slowly tend to the supercompacton limit for very large values of $\epsilon$. We have also reviewed the beautiful result of the critically coupled magnetic Skyrmion of Barton-Singer-Ross-Schroers [19], where an analytic solution is available, hence providing many studies with a clear advantage. Their solution is known to be an exact solution for any value of the coupling $\epsilon$ [19], but we find here that dropping the boundary term that leads to the correct lump size for this solution also changes the energy from the lump energy for all values of $\ell$ to the energy functional $4\pi(1-\epsilon)$. This energy functional changes sign at $\epsilon=1$ suggesting $\epsilon=1$ being the critical coupling for having the Skyrmionic phase.
@@ Acknowledgments
We thank Bruno Barton-Singer for comments on our draft. The work of S. B. is supported by the INFN special research project grant “GAST” (Gauge and String Theories). S. B. G. thanks the Outstanding Talent Program of Henan University and the Ministry of Education of Henan Province for partial support. The work of S. B. G. is supported by the National Natural Science Foundation of China (Grant No. 12071111) and by the Ministry of Science and Technology of China (Grant No. G2022026021L). R. M. thanks the Department of Physics “E. Fermi” of University of Pisa where this work was initiated.

### v3

@@ 2.2 $\operatorname{O}(3)$ sigma model limit
with $a$ an arbitrary scale of the soliton (size modulus of the $\operatorname{O}(3)$ sigma model lump). We then evaluate the energy functional for small but finite values of $\epsilon$. Since it was shown in Ref. [20] that $\widetilde{E}$ of Eq. (2.7) is the energy functional leading to a well-defined variational problem, we will plug in the above lump solution to that energy functional:
$\widetilde{E}(a)=4\pi-8\pi\kappa a+\frac{2^{p}}{p-1}h\pi a^{2},$ (2.44)
@@
For $p=1$ the solution for the lump (2.43) evaluated on $E_{0}$ diverges, see Eq. (2.44). Introducing a cutoff ($\int_{0}^{\infty}\mathop{}\!\mathrm{d}r\longrightarrow\int_{0}^{\Lambda}\mathop{}\!\mathrm{d}r$) gives
$\widetilde{E}(a)\simeq 4\pi-8\pi\kappa a+2\pi a^{2}h\log\left(1+\frac{\Lambda^{2}}{a^{2}}\right).$ (2.47)
@@
Unfortunately, it is not possible to remove the infinity leaving behind a positive results, so this estimate should be considered unphysical.
@@ 2.3 Restricted model limit
In the case of $\ell$ fixed and $\epsilon\to\infty$ we expect that the stable Skyrmion is essentially given by a stabilization between $E_{1}$ and $E_{0}$, while $E_{2}$ is negligible with respect to the other two terms. Thus the energy functional takes the simplified form444 One may wonder why we restrict the potential to the form $V_{p}$ of Eq. (2.5) and hence do not consider e.g. $V\propto(1-(n^{3})^{2})$. Although such easy-axis potential is often utilized in magnetic systems, they do not allow for the large-$\epsilon$ limit because the ground state becomes inhomogeneous [11, 52] and requires a different analysis than carried out here.
@@ 2.4 Critical coupling
We note that the energy $\widetilde{E}$ of Eq. (2.7), which has the vorticity subtracted off, leads to a different Bogomol’nyi bound [20]:
$\widetilde{E}\geq-4\pi.$ (2.66)
@@ 2.5 Analytic solution at any coupling
We note that the Bogomol’nyi completion requires us to select the critical coupling $h=\frac{\kappa^{2}}{2}$ enabling us to complete the square in the energy functional. This choice of coupling corresponds to $\epsilon=2$. However, another interesting fact about the Zeeman squared potential ($p=2$) is that inserting the Ansatz $\vartheta=2\arctan\tfrac{a}{r}$ into the full second-order Euler-Lagrange equation gives [50, 19, 21]
@@
which beautifully has the solution $a=\frac{\kappa}{h}=\ell$. This size does not contradict the BPS solution, since the first-order BPS equation required $h=\frac{\kappa^{2}}{2}$ which in the full second-order solution corresponds to the lump size $a=\frac{\kappa}{h}=\frac{2}{\kappa}$. It is interesting, however to insert the exact solution into the energy functional (2.36), which gives [19, 20]
@@
The $p=2$ case is thus applicable to many more physical systems, than just the critically coupled one [50, 19, 21]. Having the analytic solution in hand may be advantageous for further studies, including interaction and scattering etc.
Using the energy functional $\widetilde{E}$ of Eq. (2.7), where the vorticity is subtracted off, the energy becomes instead
$\widetilde{E}$ $=2\pi\int r\mathop{}\!\mathrm{d}r\;\frac{4h\kappa^{2}(h-\kappa^{2})}{(h^{2}r^{2}+\kappa^{2})^{2}}$
$=4\pi(1-\epsilon).$ (2.69)
@@ 2.6 Numerical solutions
For the general theory, we can solve equation (2.39) using a standard shooting method, beginning from the origin with the initial condition $\vartheta(0)=\pi$. The form of the solution close to the origin is $\vartheta(r)\simeq\pi-\gamma r$. By varying the angular coefficient $\gamma$, we find only one solution in the topological sector $Q=-1$ as already shown in the literature [16, 18, 20, 19, 51, 23, 21].
@@
Figure 5: Topological charge density, $\mathcal{Q}$ for (a) $p=\frac{3}{2}$ and (b) $p=1$ with $\epsilon=0.001$, $0.01$, $0.1$, $1$, $10$, $100$, $1000$. The topological charge densities of the lumps and restricted Skyrmions are shown with dashed black and red curves, respectively. In the $p=1$ case, the size of the lump is selected for illustrative purpose, corresponding to the solutions in Fig. 4.
@@
Figure 6: Energy density: (a) $p=\frac{3}{2}$ (b) $p=1$ for $\epsilon=0.001,0.01,0.1,1$ (solid blue lines) and with the boundary term (2.9) subtracted off (dashed dark-blue lines). Energy density: (c) $p=\frac{3}{2}$ (d) $p=1$ for $\epsilon=1,10,100,1000$ and finally (e) $p=\frac{3}{2}$ (f) $p=1$ and $\epsilon=1,10,100,1000$ with the boundary term (2.9) subtracted off.
Fig. 6 shows the energy densities for the sigma-model lump limit ($\epsilon\to 0$) in panels (a) and (b) with and without the boundary term for $p=\frac{3}{2}$ and $p=1$, respectively, and for the restricted limit ($\epsilon\to\infty$) in panels (c) and (d) for $p=\frac{3}{2}$ and $p=1$, respectively, in panels (e) and (f) $p=\frac{3}{2}$ and $p=1$, respectively, with the boundary term (2.9) subtracted off. In both the charge density as well as in the energy density, the discontinuity that develops in the $p=1$ case for $\epsilon\to\infty$ is clearly seen in Figs. 5(b) and 6(d), respectively. We note that the energy density is divergent (singular) in the $p=\frac{3}{2}$ case at $\rho=1$ when the boundary term is taken into account, see Fig. 6(c), but is convergent (finite) when the boundary term is dropped, see Fig. 6(e). We also note that the total energy (the integral) is unchanged in the $p=1$ case (see Figs. 6(d) and 6(f)), but the energy density is slightly modified by the presence of the boundary term.
Figure 7: Energy in the lump limit ($\epsilon\to 0$) as a function of $\epsilon$ for $p=\tfrac{3}{2}$ and $p=1$. (a) The blue dashed line corresponds to the analytical energy formula (2.46) for the lump in the limit of small but nonvanishing $\epsilon$ in the case of $p=\frac{3}{2}$. The red dashed line instead is the energy of Eq. (2.50) with the lump size taken from Fig. 4. (b) The energy shown up to larger values of $\epsilon$ revealing the sign change near $21<\epsilon<$ 2. The case $p=\frac{3}{2}$ is shown for the full energy (solid blue line) and with the boundary term (2.9) subtracted off (dashed blue line).
We show the total energy of the numerical solutions in the lump limit ($\epsilon\to 0$) for the $p=\frac{3}{2},1$ cases in Fig. 7(a). We notice that the $p=\frac{3}{2}$ case is well described by the energy functional (2.46) with the correct lump size (2.45). This confirms that discarding the boundary term (2.9) is appropriate for computing the energies in the lump limit, which is also the case where the variational problem is well defined [20]. In the $p=1$ case, we show the regularized energy (2.47) with a dashed red curve and lump size $a$ as chosen in Fig. 4. We show the sign change of the energy in the $p=\frac{3}{2},1$ case in Fig. 7(b) which appears to take place around $21<\epsilon<$ 2, which is consistent with the lump prediction (2.46) which for $p=\frac{3}{2}$ should take place at $\epsilon=\epsilon^{\rm crit}=\sqrt{2}$.
@@ 2.7 Comparison with experiments
We have confirmed that the magnetic Skyrmion model, up to rescaling of energy and length, is essentially a one-parameter model (2.27)-(2.28). This parameter, $\epsilon$, controls the relation between $E_{2}$ and $E_{1}+E_{0}$. In both the large-$\epsilon$ and the small-$\epsilon$ limit, analytic approximations are available, see Secs. 2.2 and 2.3. It may be interesting to explore various phenomenological realizations of chiral magnets and determine the order of magnitude of the parameter $\epsilon$ that can be realized.
@@
where $\operatorname{J}$ is the unit Joule, not to be confused with the interaction strength of the Heisenberg Hamiltonian. Note that in general $\ell$ can change significantly, viz. one order of magnitude. For instance, if the we fix $\ell=100\operatorname{nm}$ and $\lambda\in\mathcal{O}(10^{1})$ nm as before, we find that $\epsilon$ changes one order of magnitude, i.e. $\epsilon\simeq 100$. For materials other than MnSi, e.g. Fe1-xCoxSi or Mn1-xFexGe, $\lambda$ can also be of the order $\mathcal{O}(10^{2})\operatorname{nm}$. Therefore, considering the other two cases where $\lambda$ is fixed as $\lambda\sim 10^{2}\operatorname{nm}$ and $\ell$ takes its maximum and minimum values, we find the relevant ranges for $\epsilon$:
@@
In Secs. 2.4 and 2.5, we saw that in the special physical case of $p=2$, which describes chiral magnets with anisotropy interactions, any spherical magnetic Skyrmion can be represented by the analytical lump solution, regardless of its DM coupling, $\kappa$. However, the potential $V_{p=2}$ corresponds to the sum of a magnetic field and anisotropy (easy-plane) contributions. To realize this potential, the couplings of the anisotropy term must be exactly half of that of the Zeeman term. It turns out to be possible [15]. Furthermore, since the critical coupling regime corresponds to the condition $\forall\leavevmode\nobreak\ \ell,\ \epsilon=2$, which satisfies the inequalities (2.73), we suggest that, in principle, a BPS magnetic Skyrmion could be realized experimentally.
@@ 3 Conclusion
In this paper, we have studied magnetic Skyrmions in chiral magnets with three different potentials, corresponding to $p=1$, $p=\frac{3}{2}$ and $p=2$ in the limits of the control parameter $\epsilon=\frac{\kappa^{2}}{h}=\frac{4D^{2}}{|B|J}$ becoming very large and going to zero. The $p=1$ case is simply the Zeeman potential term and the $p=2$ case is a particular combination of the easy-plane (anisotropy) and the Zeeman potentials. The $p=\frac{3}{2}$ case is probably somewhat more academic. The $\epsilon$ vanishing limit reduces the Skyrmion to a sigma model lump with a lump size that we accurately estimated, except for the case of the Zeeman term ($p=1$). The large-$\epsilon$ limit instead corresponds to the “restricted” magnetic Skyrmions, for which we found the exact analytic solutions as well. In the $p=\frac{3}{2}$ case, the restricted solution is a compacton, similar in nature to those found in the baby-Skyrme [43, 44, 45] and 3D Skyrme models [46]. However, in the $p=1$ case corresponding to the Zeeman term, we find a new type of restricted solution, which we dub a supercompacton, since it becomes discontinuous. We provide numerical solutions showing that they slowly tend to the supercompacton limit for very large values of $\epsilon$. We have also reviewed the beautiful result of the critically coupled magnetic Skyrmion of Barton-Singer-Ross-Schroers [19], where an analytic solution is available, hence providing many studies with a clear advantage. Their solution is known to be an exact solution for any value of the coupling $\epsilon$ [50, 19, 21], but we find here that dropping the boundary term that leads to the correct lump size for this solution also changes the energy from the lump energy for all values of $\ell$ to the energy functional $4\pi(1-\epsilon)$. This energy functional changes sign at $\epsilon=1$ suggesting $\epsilon=1$ being the critical coupling for having the Skyrmionic phase. We finally emphasize that our study, within the defined restricted limit, provides analytical solutions for magnetic Skyrmions in certain magnetic material, as shown in Sec. 2.7.
@@ Acknowledgments
We thank Bruno Barton-Singer and Calum Ross for comments on our draft. The work of S. B. is supported by the INFN special research project grant “GAST” (Gauge and String Theories). S. B. G. thanks the Outstanding Talent Program of Henan University and the Ministry of Education of Henan Province for partial support. The work of S. B. G. is supported by the National Natural Science Foundation of China (Grant No. 12071111) and by the Ministry of Science and Technology of China (Grant No. G2022026021L). R. M. thanks the Department of Physics “E. Fermi” of University of Pisa where this work was initiated.